Can you please help?
.
Please solve  64 = 4^(2x) 16^(3x^2) , it is 16 to the exponent of (3x^2)
.
Thanks ahead of time, Levi
: Can you please help?
.
Please solve  64 = 4^(2x) 16^(3x^2) , it is 16 to the exponent of (3x^2)
.
Thanks ahead of time, Levi

Answer by Alan3354(1926)

Please solve  64 = 4^(2x) 16^(3x^2) , it is 16 to the exponent of (3x^2)
----------------
 64 = 4^(2x) 16^(3x^2)
 2^6 = 2^(4x) * 2^(12x^2)
 2^6 = 2^(4x + 12x^2)
12x^2 + 4x = 6
6x^2 + 2x - 3 = 0
x = -1/6 + sqrt(19)/6
x = -1/6 - sqrt(19)/6